# MONKEY: Masking ON KEY-Value Activation Adapter for Personalization

### v2

@@ 1 Introduction
Diffusion models have become a dominant paradigm for image generation (ho2020denoisingdiffusionprobabilisticmodels; rombach2022highresolutionimagesynthesislatent), competitive with autoregressive models (parmar2018imagetransformer; xiong2025autoregressivemodelsvisionsurvey), and largely superseding GANs (goodfellow2014generativeadversarialnetworks). However, despite their expressive power, end users often desire more control over who or what appears in the generated image. Personalization methods address this need by allowing a model to incorporate a specific person or object into new generations while retaining the flexibility of text-based conditioning and adhering to the text prompt.
Existing personalization methods fall into the categories of test-time fine tuning, such as Dreambooth (ruiz2023dreamboothfinetuningtexttoimage) and Textual Inversion (gal2022imageworthwordpersonalizing), or adapter-based methods like IP-Adapter (ye2023ipadaptertextcompatibleimage) and InstantID (wang2024instantidzeroshotidentitypreservinggeneration). Both categories often struggle to balance fidelity to the subject with faithfulness to the text prompt. In particular, when strong visual features of the subject dominate, the model tends to reproduce the original subject image rather than composing it naturally into new scenes or contexts. This results in limited generalization and weaker prompt alignment.
@@ 2.2 Personalization
Text-to-image diffusion models Croitoru_2023; chen2025comprehensive have been incredibly impressive in their ability to generate diverse, realistic images. While generally prompted on just text, a diffusion model can also be prompted on a particular instance of a category. For example, a user may want to generate pictures of a specific cat or person, instead of a generic cat or person. This is the task of Personalization. A user should be able to supply a few, or even a single image of a concept (such as a face, object or person) and be able to reliably generate new images of that concept that match input prompts. We follow the dichotomy of (zhang2024survey) of dividing personalization methods into test-time fine tuning or pretrained adaptation. The former group requires the diffusion model to be retrained on each new concept. Some examples are Dreambooth (ruiz2023dreamboothfinetuningtexttoimage), Textual Inversion (gal2022imageworthwordpersonalizing) and Imagic (kawar2023imagictextbasedrealimage). Pretrained adaptation does not require any new training. Concept images can be supplied to the diffusion model and personalized images immediately created at inference. Some examples include ID-Aligner (chen2024idalignerenhancingidentitypreservingtexttoimage), Face2Diffusion (shiohara2024face2diffusionfasteditableface), and most relevant to this work, IP-Adapter ye2023ipadaptertextcompatibleimage. IP-Adapter embeds the concept image using CLIP (radford2021learningtransferablevisualmodels) and learns an embedding for each layer of the UNet. The layer-wise embedding attends to the output of each layer using cross-attention.
@@ 2.3 Extracting Features from UNets
Unets (ronneberger2015unetconvolutionalnetworksbiomedical) are a common backbone model for diffusion (ho2020denoisingdiffusionprobabilisticmodels), where they are trained to predict the noise from a corrupted image. At inference time, they gradually reduce pure noise to an actual image. The later layers of models like ResNet (he2015deepresiduallearningimage) or VGG (simonyan2015deepconvolutionalnetworkslargescale), can be used as semantic feature extractors, while earlier layers extract stylistic information. For UNets, the center "bottleneck" layers extract semantic content features, while the earlier layers learn style features (haas2024discoveringinterpretabledirectionssemantic; xu2024freetunersubjectstyletrainingfree; frenkel2024implicitstylecontentseparationusing; schaerf2025trainingfreestylecontenttransfer). This property has been used for personalization, such as doing text inversion using a new token for each layer for each new concept (voynov2023pextendedtextualconditioning; agarwal2023imageworthmultiplewords). Different timesteps of the inference process are also more influential than others in determining the content features of the final images (yu2023freedomtrainingfreeenergyguidedconditional; zhang2023prospectpromptspectrumattributeaware; agarwal2023imageworthmultiplewords).
@@ 3.1 Attention Maps
The original IP-Adapter ye2023ipadaptertextcompatibleimage checkpoint, for each layer of the diffusion model, produces four tokens that are concatenated to the text tokens during the cross attention stage. During cross attention, the intermediate features $\phi_{\ell}$ are passed to each layer, and multiplied by the query matrix to get $Q$ and then the text and image token embeddings are multiplied by the key and value matrices to get $K$ and $V$. The final output of the attention layer is $\textrm{softmax}(\frac{QK^{T}}{\sqrt{d}})\cdot V$. The $\textrm{softmax}(\frac{QK^{T}}{\sqrt{d}})$ component represents the degree of correspondence between the semantic embeddings of each token and each spatial location in the image. Given that deeper levels of UNets correspond to semantic features, we decided to plot the activations of the text and ip tokens of an intermediate layer of each part of the generated images.
@@
We use the https://huggingface.co/SimianLuo/LCM_Dreamshaper_v7 checkpoint from huggingface, which is a latent consistency (luo2023latentconsistencymodelssynthesizing) distillation of https://huggingface.co/Lykon/dreamshaper-7. For the IP-Adapter, we used the standard https://huggingface.co/h94/IP-Adapter checkpoint. We used the second transformer in the first up-block layer named up_blocks.1.attentions.1.transformer_blocks.0.attn2. For some visualization of the attention maps of other layers, refer to A.2
@@ 4.1 Models
We compared our method to a few other test-time personalization methods. Given that our method did not require training any new models or weights, we chose works that also did not train any new weights, those being FreeGraftor (yao2025freegraftortrainingfreecrossimagefeature), RectifID (sun2024rectifidpersonalizingrectifiedflow), MASA (cao2023masactrltuningfreemutualselfattention) and TF-I2I (hsiao2025tfti2itrainingfreetextandimagetoimagegeneration), in addition to the baseline ip-adapter checkpoint with the ip token scale set to 0.5 and 1.0.
@@ 4.2 Data
We used two image datasets: the Dreambooth dataset (ruiz2023dreamboothfinetuningtexttoimage), consisting of objects and animals, and our own curated dataset of illustrations of characters from the card game Magic the Gathering, which featured mostly humanoids with various fantastical features and diverse appearances. For the text prompts, we used a set of prompts relating to background and location such as on top of green grass with sunflowers around it and on a cobblestone street, generated by asking ChatGPT (openai2023gpt4). A complete list can be found in A.1.
@@ 4.3.1 Automated Evaluation
We used DINOv2 Identity (oquab2024dinov2learningrobustvisual), CLIP Image and CLIP Text embedding similarity. The first two measure similarity between the source and generated images, and the third measures similarity between the text prompt and generated image. Scores for each dataset are shown in Table 1. We provide visual comparisons using some handpicked results in Table 2
@@ 5 Conclusion and Future Work
We presented a simple inference-time masking strategy that improves the balance between subject image alignment and text prompt alignment in adapter-based personalization of diffusion models. By leveraging the implicit subject masks already present in IP-Adapter activations, our method enhances compositional control without training any new weights or retraining pre-existing weights. Future work can extend this approach toward multi-subject personalization and combining MONKEY Adapter with complementary personalization techniques such as ControlNet (zhang2023addingconditionalcontroltexttoimage).

### v3

@@ 1 Introduction
Diffusion models have become a dominant paradigm for image generation (Ho et al., 2020; Rombach et al., 2022), competitive with autoregressive models (Parmar et al., 2018; Xiong et al., 2025), and largely superseding GANs (Goodfellow et al., 2014). However, despite their expressive power, end users often desire more control over who or what appears in the generated image. Personalization methods address this need by allowing a model to incorporate a specific person or object into new generations while retaining the flexibility of text-based conditioning and adhering to the text prompt.
Existing personalization methods fall into the categories of test-time fine tuning, such as Dreambooth (Ruiz et al., 2023) and Textual Inversion (Gal et al., 2022), or adapter-based methods like IP-Adapter (Ye et al., 2023) and InstantID (Wang et al., 2024). Both categories often struggle to balance fidelity to the subject with faithfulness to the text prompt. In particular, when strong visual features of the subject dominate, the model tends to reproduce the original subject image rather than composing it naturally into new scenes or contexts. This results in limited generalization and weaker prompt alignment.
@@ 2.1 Personalization
Text-to-image diffusion models Croitoru et al. (2023); Chen et al. (2025) have been incredibly impressive in their ability to generate diverse, realistic images. While generally prompted on just text, a diffusion model can also be prompted on a particular instance of a category. For example, a user may want to generate pictures of a specific cat or person, instead of a generic cat or person. This is the task of Personalization. A user should be able to supply a few, or even a single image of a concept (such as a face, object or person) and be able to reliably generate new images of that concept that match input prompts. We follow the dichotomy of (Zhang et al., 2024) of dividing personalization methods into test-time fine tuning or pretrained adaptation. The former group requires the diffusion model to be retrained on each new concept. Some examples are Dreambooth (Ruiz et al., 2023), Textual Inversion (Gal et al., 2022) and Imagic (Kawar et al., 2023). Pretrained adaptation does not require any new training. Concept images can be supplied to the diffusion model and personalized images immediately created at inference. Some examples include ID-Aligner (Chen et al., 2024), Face2Diffusion (Shiohara and Yamasaki, 2024), and most relevant to this work, IP-Adapter Ye et al. (2023). IP-Adapter embeds the concept image using CLIP (Radford et al., 2021) and learns an embedding for each layer of the UNet. The layer-wise embedding attends to the output of each layer using cross-attention.
@@ 2.2 Extracting Features from UNets
Unets (Ronneberger et al., 2015) are a common backbone model for diffusion (Ho et al., 2020), where they are trained to predict the noise from a corrupted image. At inference time, they gradually reduce pure noise to an actual image. The later layers of models like ResNet (He et al., 2015) or VGG (Simonyan and Zisserman, 2015), can be used as semantic feature extractors, while earlier layers extract stylistic information. For UNets, the center "bottleneck" layers extract semantic content features, while the earlier layers learn style features (Haas et al., 2024; Xu et al., 2024; Frenkel et al., 2024; Schaerf et al., 2025). This property has been used for personalization, such as doing text inversion using a new token for each layer for each new concept (Voynov et al., 2023; Agarwal et al., 2023). Different timesteps of the inference process are also more influential than others in determining the content features of the final images (Yu et al., 2023; Zhang et al., 2023b; Agarwal et al., 2023).
@@ 3.1 Attention Maps
The original IP-Adapter Ye et al. (2023) checkpoint, for each layer of the diffusion model, produces four tokens that are concatenated to the text tokens during the cross attention stage. During cross attention, the intermediate features $\phi_{\ell}$ are passed to each layer, and multiplied by the query matrix to get $Q$ and then the text and image token embeddings are multiplied by the key and value matrices to get $K$ and $V$. The final output of the attention layer is $\textrm{softmax}(\frac{QK^{T}}{\sqrt{d}})\cdot V$. The $\textrm{softmax}(\frac{QK^{T}}{\sqrt{d}})$ component represents the degree of correspondence between the semantic embeddings of each token and each spatial location in the image. Given that deeper levels of UNets correspond to semantic features, we decided to plot the activations of the text and ip tokens of an intermediate layer of each part of the generated images.
@@
We use the https://huggingface.co/SimianLuo/LCM_Dreamshaper_v7 checkpoint from huggingface, which is a latent consistency (Luo et al., 2023) distillation of https://huggingface.co/Lykon/dreamshaper-7. For the IP-Adapter, we used the standard https://huggingface.co/h94/IP-Adapter checkpoint. We used the second transformer in the first up-block layer named up_blocks.1.attentions.1.transformer_blocks.0.attn2. For some visualization of the attention maps of other layers, refer to A.2
@@ 4.1 Models
We compared our method to a few other test-time personalization methods. Given that our method did not require training any new models or weights, we chose works that also did not train any new weights, those being FreeGraftor (Yao et al., 2025), RectifID (Sun et al., 2024), MASA (Cao et al., 2023) and TF-I2I (Hsiao et al., 2025), in addition to the baseline ip-adapter checkpoint with the ip token scale set to 0.5 and 1.0.
@@ 4.2 Data
We used two image datasets: the Dreambooth dataset (Ruiz et al., 2023), consisting of objects and animals, and our own curated dataset of illustrations of characters from the card game Magic the Gathering, which featured mostly humanoids with various fantastical features and diverse appearances. For the text prompts, we used a set of prompts relating to background and location such as on top of green grass with sunflowers around it and on a cobblestone street, generated by asking ChatGPT (OpenAI, 2023). A complete list can be found in A.1.
@@ 4.3.1 Automated Evaluation
We used DINOv2 Identity (Oquab et al., 2024), CLIP Image and CLIP Text embedding similarity. The first two measure similarity between the source and generated images, and the third measures similarity between the text prompt and generated image. Scores for each dataset are shown in Table 1. We provide visual comparisons using some handpicked results in Table 2
@@ 5 Conclusion and Future Work
We presented a simple inference-time masking strategy that improves the balance between subject image alignment and text prompt alignment in adapter-based personalization of diffusion models. By leveraging the implicit subject masks already present in IP-Adapter activations, our method enhances compositional control without training any new weights or retraining pre-existing weights. Future work can extend this approach toward multi-subject personalization and combining MONKEY Adapter with complementary personalization techniques such as ControlNet (Zhang et al., 2023a).